Can someone help, I keep applying the formula, but get wrong answer.
Use A= P(1+rt) to find t, when
A=$966
P=$840
r= 5%
Write out steps so I can see what I'm doing wrong, thankyou.

Found 2 solutions by unlockmath, sachi:
Answer by unlockmath(1688)

Hello,
Here we go:A= P(1+rt)
966=840(1+.05t)
Divide by 840:
1.15=1 +.05t
Subtract 1
.15 = .05t
Divide by .05 to get:
t = 3 years
Make sense?
RJ
www.math-unlock.com

Answer by sachi(548)

Use A= P(1+rt) to find t, when
A=$966
P=$840
r= 5%=5/100=.05
A= P(1+rt)
or 1+rt=A/P
or rt=A/P-1=966/840-1=23/20-1=3/20
or t=3/r*20=3/0.05*20=3/1.00=3 years